A solution is prepared by adding 5g of a substance x to 18 gm water. calculate the mass percent of the solute
1
Expert's answer
2020-11-10T14:01:56-0500

"mass\%_{solute}=\frac{m_{solute}}{m_{solute}+m_{solvent}}=\frac{5g}{5g+18g}\times100\%=21.7\%"


Answer: 21.7%